Sales Mix and Break-Even Analysis

Megan Company has fixed costs of $402,380. The unit selling price, variable cost per unit, and contribution margin per unit for the company's two products follow:

Product Selling Price Variable Cost per Unit Contribution Margin per Unit
QQ $140 $60 $80
ZZ 180 140 40

The sales mix for Products QQ and ZZ is 55% and 45%, respectively. Determine the break-even point in units of QQ and ZZ. If required, round your answers to the nearest whole number.

a. Product QQ  units

b. Product ZZ  units
Answer #1
Fixed costs 402380
/ Weighted average Contribution Margin per Unit 62 =(80*55%)+(40*45%)
Overall break even point 6490
a. Product QQ units 3570 or 3569 =6490*55%
b. Product ZZ units 2921 or 2920 =6490*45%